OpenStudy (anonymous):

would 81x^2+72x+16 be (9x+9)(9x+7) in factored form?

OpenStudy (anonymous):

no because \[7\times 9\neq16\]

OpenStudy (anonymous):

so 8 and 2?

OpenStudy (anonymous):

but what about the 72?

OpenStudy (anonymous):

no. you got the 9 right. what is 16 the square of?

OpenStudy (anonymous):

that is the one that will give it to you

OpenStudy (anonymous):

16^2 is 256

OpenStudy (anonymous):

question is "what is 16 the square of" or "what number squared gives 16"

OpenStudy (anonymous):

4

OpenStudy (anonymous):

yes and that completes the problem (and completes the square)

OpenStudy (anonymous):

(9x+4)^2

OpenStudy (anonymous):

the end part would be 63 instead of 16

OpenStudy (anonymous):

So (9x+4)(9x+4) is the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

gblover multiply out \[(9x+4)^2\] and see that it is right

OpenStudy (anonymous):

9x*9x=81x^2, 9x*4*2=72x, 4*4=16

OpenStudy (anonymous):

first term will be \[81x^2\] last term will be \[4^2=16\] and middle term will be \[2\times 9\times 4x\]

OpenStudy (anonymous):

what pmel11 said